Report for April 2018

Report Date
April 2018

Section One: Programme Commentary, RAG and Resources Update

Programme Commentary

In this reporting period work has continued on all active projects. Good progress has been made with the Grouper Upgrade, with UAT completed and deployment now scheduled. The BI Annual Upgrade remains RED as due to ongoing performance issues on TEST. The performance issues on DEV have been resolved. There has been a change of PM on two projects with the departure of Nick Graham.

 

RAG Status and Forward Look

 

Resources

 

 Section Two: Projects Not Started

Project Info     Project Manager's Commentary
COM040 JIRA Upgrade  Scheduled to start in May 2018

 

Section Three: Projects in Progress

 

Project Info Project Manager's Commentary

COM036 BI - Business Objects Upgrade

Compliance

Milestones

11-May-2018 Test build complete
25-May-2018 Integration Testing sign off
29-Jun-2018 Acceptance Testing sign off
14-Jul-2018 Deploy to LIVE
27-Jul-2018 Deployment Sign Off Review
03-Aug-2018 Closure

RAG Status 

RED

 

Report by Chris Konczak for April 2018

Achievements 

  • Itelligence completed functionality testing. 1 fail. This is due to BI service accounts not having correct permissions assigned. This has been assigned to Dev Tech to resolve.
  • Itelligence availability for deployment in July 2018 secured for out of hours

Issues

  • Performance issues not resolved. Production are continuing investigating and carrying out testing on the hosting infrastructure to identify where the issue is. Changes will be made to the infrastructure to test resolution scenarios.
  • Handover of environments - conditions not yet met. Itelligence to re-test the Oracle performance problem on 3 environments and feedback to Production. Once this is satisfactory then
  • WIS requested that the project is to continue to report a red RAG until the performance technical issue has been resolved.

Next Steps 

  • Itelligence to provide effort estimates for applying a patch that should resolve the security migration bug. This will allow us to migrate data using the wizard and not the command line
  • Milestones to be replanned (environment builds and integration sign off). Acceptance sign off, Deployment to LIVE and closure milestones are to remain as they are.  
  • DEV server configuration, content migration and sign off to be achieved
  • TEST server configuration and data migration to commence.

RAG status is RED - Issue: Performance problems on VM servers (Dev and TEST) have not been resolved. Delivery: Environment handover from Itelligence has been delayed, relating partly but not solely to this issue. No current impact on deployment date, scheduled for July. IS Apps Production Management are carrying out testing and investigation on different hosts week commencing 30/04/2018, by running on fast and slower hosts and recording times. If +/- 30% issue will be closed, otherwise escalated to ITI. No SMT action needed.

 

COM037 - Grouper Upgrade

Compliance

Milestones

Target Datesort descending Title
08-May-2018 Deliver
15-May-2018 Deployment Review
29-May-2018 Closure Review

RAG Status 

GREEN

Report by David Watters for April 2018

Achievements 

  • UAT carried out successfully.
  • Delivery re-planned, approved by stakeholders and assured by WIS. 
  • Deployment checklist and planning prepared. 

Issues

  • None.

Next Steps 

  • Cutover and delivery.  

The project is reporting GREEN.

COM038 IDM SOA Replacement

Compliance

Milestones

 

20-Apr-2018    Acceptance Review

25-May-2018    Deliver (Final Report)

15-Jun-2018    Closure

RAG Status

AMBER

Report by Adam Wadee for April 2018

Achievements

  • During the PoC demo, the Project Team demonstrated that they successfully installed and configured the WSO2 Integrator, and used it to integrate with our existing IDM SOA components.  This was seen as good progress in terms of looking for a replacement for Oracle SOA ESB (or Oracle SOA Mediator as it’s now called).  However it was noted at the meeting that if we did replace Oracle SOA Mediator with the WSO2 Integrator, this would only be a first, small step towards replacing Oracle SOA completely.  To completely remove Oracle SOA we’d also need to replace the BPEL processes, which would need significant development.  As the future strategy for SOA and IDM is being reviewed out-with this project, it has been recommended that we upgrade the SOA infrastructure to make it supportable for the foreseeable future (to 12c).  As agreed with colleagues from Dev Services and Production Management, Dev Tech are currently investigating the requirements for upgrading SOA on Dev.  A review of the estimates and tasks involved with the upgrade will be done and if we have sufficient budget available in COM038 we’ll progress with the upgrade on Dev.

Issues

  • As this is a change from the current project scope, approval from the project team will be obtained to update the scope and milestones once we have information about the 12c upgrade. 

Next steps

  • Review the estimates for upgrading to SOA 12c.  If it is agreed to progress this will result in a change of scope for the project and revised milestones.

As the next steps for the project have not yet been confirmed, the project will be reported as AMBER this month.

COM039 Wiki Upgrade

Compliance

Milestones

14-May-2018 Wiki Integration signoff (excl EdGEL)
08-Jun-2018 Wiki Acceptance sign off
20-Jun-2018 Wiki Deploy to LIVE (excl EdGEL)
26-Jun-2018 DSOR
13-Jul-2018 Project Closure

RAG Status 

GREEN

 

Report by Chris Konczak for April 2018

Achievements in Last Period

  • Collaborative Editing fix on DEV and TEST completed
  • TEST database peer testing completed
  • TEST database sign off achieved
  • New version of Confluence installed on TEST completed
  • Peer testing of Confluence completed on DEV and TEST completed
  • Database upgrade on LIVE achieved

Issues

  • No Issues.

Next Steps

  • Load testing
  • Wiki Integration sign off
  • Wiki UAT to commence
  • Database deployment sign off review

COM041 Decommission legacy Windows hosting

Compliance

 

Milestones

Target Datesort descending Title
02-Jul-2018 Migration complete
13-Jul-2018 Closure complete

 

RAG Status

GREEN

 Report  by Megan Graham for April 2018

 Achievements This Period

  • Web migrations are continuing to the new infrastructure. Many of them are complicated and time consuming, requiring additional time from Service manager to support the technical resources, but we are still within the project tolerance for timescales.
  • Wiki has been updated with the process and plan for migrations.
  • Recruitment extension and interview of Student intern resource.

 Issues

  • Recruitment exercise has been unsuccessful and has led to an extension of the time period, in the hope that a suitable candidate will be found.  Contingency is for the Project services resource to extend their work to perform the less complex migrations, with support and training from Service Management.

 Next Steps

  • Make headway towards completing the migrations to the new infrastructure.

COM042 GDPR Preparation for Applications

Milestones

Target Datesort descending Title         

11-May-2018

Acceptance of Action Plan and Final Report

18-May-2018

Project Closure

 

RAG Status

GREEN

Report by Emma McNab for April 2018

Achievements

  • Finalization of Internal Register 
  • Action plan template
  • Project Manager hand over

Issues 

n/a

Next Steps

  • Sign Off of Internal Register
  • Privacy Notices completion
  • Report on exceptions and compliance actions to be completed

 

Section Four: Closed or Withdrawn Projects

Project Info Project Managers Commentary

COM034

Annual Wiki Upgrade

Compliance

Report by Chris Konczak

This project was closed on 08/08/17  (WIS Notes 09/09/17)

Please see the Completion Report for further information

COM035 

JIRA Upgrade 2017

Compliance

Report by Tim Gray

This project was closed on 14/08/17  (WIS Notes 11/08/17)

Please see the Completion Report for further information

COM032

Rollout Academic Timetable Integration with Office 365 Calendar

Discretionary

Report by Chris Copner

This project was closed on 18/08/17  (WIS Notes 18/08/17)

Please see the Completion Report for further information

COM031

Serengeti Document Management Server Upgrade

Compliance

Report by Morna Findlay

This project was closed on (WIS Notes 23/03/18)

Please see the Completion Report for further information.

 

Section Five: Actions/Notes from monthly meeting held on 15/05/2018

Project Action / Notes Responsibility    Status
COM036

The migration tool is not fit for purpose. SC to await feedback from vendor today if not successful then escalate through AC

PJ to speak to Mark McG re figures on the technical issues. PJ suggested stress test on LIVE with 5/6 users - SC to speak to AMcF to organise.

CK to progress weekend work - has everyone responded? Have ITI confirmed their availability for DNS change? Project team to discuss if weekend work is really necessary?

SC/AC

PJ/SC/AMcF

TG/CK

 
COM038

Project should aim to complete as much infrastructure work this year as possible to utilise the available resources. 

AW

 
COM039 License issue to be clarified - TG follow up with CK/SK TG/CK  
COM040 JIRA Upgrade project to be moved to start in June. Forward Look to be updated and PRICCL raised. DW  
AttachmentSize
Image icon fl-apr18-2.png43.72 KB